OpenStudy (anonymous):

Which of the following is the rational exponent expression of ^3√4n?

myininaya (myininaya):

\[\sqrt[3]{4n}=(4n)^\frac{1}{3}\] is this what you are looking for?

OpenStudy (anonymous):

Yes, Thank you so much. (:

myininaya (myininaya):

\[\sqrt[a]{x^b}=x^\frac{b}{a}\]